OpenStudy (anonymous):

Phil’s age is 7 years more than times Peter’s age. Also, 4 times Phil’s age is 2 years less than twice Peter’s age. If x is Peter’s age in years, identify the equation that represents this situation and identify the solution to the equation.

OpenStudy (retireed):

Is here a missing value in your question statement ? .... Phil's age is 7 years more than ______ times Peter's age. The blank should have a number or a word like twice.

OpenStudy (retireed):

First define your variables... Ph = Phil X = Peter Y unknown multiplier Eq 1) Ph - X*Y = 7 I can only write two equation, but have 3 unknowns Eq 2) 2X - 4Ph = 2 Eq 3) Ph = 7 + XY Eq 4) 2X - 4(7+XY) = 2 2X - 28 - 4XY = 2 X ( 2 - 4Y ) = 30 X = 30 / (2 - 4Y) Y has a limited positive domain Let Y = 1/4 X = 30 Ph would = 14.5 If Peter is 30, Phil is 14.5 years old